# Daniil Orel1, Indraneil Paul2, Iryna Gurevych1,2, and Preslav Nakov1

### v1

@@ Abstract
Daniil Orel1, Indraneil Paul2, Iryna Gurevych1,2, and Preslav Nakov1 1 Mohamed bin Zayed University of Artificial Intelligence (MBZUAI), UAE 2 Ubiquitous Knowledge Processing Lab (UKP Lab) Department of Computer Science and Hessian Center for AI (hessian.AI) Technische Universität Darmstadt {name.surname}@mbzuai.ac.ae; {name.surname}@tu-darmstadt.de
@@ 2.1 AI-Generated Text Detection
Tools such as MOSS (Puryear and Sprint, 2022) have shown some effectiveness in identifying AI-generated code, since their style is out of the ordinary distributions of student solutions. However, Pan et al. (2024) and JianWang et al. (2024) have shown that detectors such as GPT-Zero often fail when applied to code rather than text. This critical observation, backed up by our experiments in Section 4, highlights the inadequacy of directly porting generic text-based models to the code domain and strongly motivates the creation of code-specific detection strategies and specialised datasets. Our work responds to this need by providing a large-scale, multifaceted suite specifically curated for AI-generated code, designed to foster the development and rigorous testing of detection techniques attuned to the unique characteristics of programming languages and AI-generated software.
@@ 3 The DroidCollection Corpus
We compare our dataset with existing ones in Table 1, which shows that our dataset is not only among the largest to date, but also captures a broader range of variations. Additional details about key characteristics are provided in Section B.3.
In this section, we detail the curation of the human-generated, machine-generated, and machine-refined splits of DroidCollection. The adversarially humanised data collection is deferred to Section 3.4.
@@ 3.1 Human-Authored Code Acquisition
To build the dataset, we collected human-written samples from multiple sources, covering C++/C, C#, Go, Java, JavaScript and Python languages. Then, we generated code, using base and instruction-tuned LMs from 11 model release families, namely: Llama (Grattafiori et al., 2024), CodeLlama, GPT-4o, Qwen (Qwen et al., 2025), IBM Granite (Mishra et al., 2024), Yi (AI et al., 2025), DeepSeek (Guo et al., 2024), Phi (Abdin et al., 2024), Gemma (Gemma et al., 2024), Mistral (AI, 2025), Starcoder (Li et al., 2023c). The list of generators per model family is detailed in Appendix A. Our dataset covers three domains: general use code, algorithmic problems, and research/data-science code.
@@ 3.2 AI-Authored Code Generation
Since the data from sources such as CodeNet and StarcoderData do not contain any instructions, we decided to apply inverse instructions to transform code from these datasets into instructions, which can be used to prompt LMs. In our case, the method of preparing inverse instructions was similar to that described in InverseCoder (Wu et al., 2025): we passed the code snippets to an LM, asking it to build a summary, and a step-by-step instruction that can be given to an LM to generate a similar code. The main difference between our approach and that of InverseCoder is that we sought to minimise the costs of the generation and did not split the summarisation and instruction generation into separate LM calls. However, in cases where a summary could be extracted from the response but the instruction could not, we used the summary to re-generate the instruction. This experiment with details about the prompts and the models we used is illustrated in Section B.1. This type of generation allows us to cover a wide range of prompts, simulating a diversity of user-LM interactions, which is common in the real world.
@@
To explore how large language models can simulate the behaviour profiles of real programmers, we took inspiration from the PersonaHub dataset (Ge et al., 2024). We first generated a diverse set of programmer profiles, and then used an LM to create programming tasks that can typically be performed by programmers of such types. These tasks, along with their corresponding descriptions, termed DroidCollection-Personas, were then used to generate code samples. More details about the DroidCollection-Personas are outlined in the Section B.2.
@@ 3.4 Adversarial Samples
To this end, we introduce two types of adversarial attacks: prompt-based attacks and preference-tuning-based attacks. In the prompt-based setting, we construct adversarial prompts by instructing the model to “write like a human” in multiple ways, relying on the models’ parametric knowledge of how to produce outputs that mimic human-authored code and thus challenge detection systems. In the preference-tuning-based setting, we curate DroidCollection-Pref, a dataset of 157K paired examples consisting of human-written and LM-generated code responses to the same prompt. Using DroidCollection-Pref, we train LMs with up to 9B parameters –including LLaMA, Qwen, and Yi–, using LoRA (Hu et al., 2022) with rank 128 and DPO for two epochs. These models’ output distributions are, in effect, steered towards preferring human-like code, making them less likely to contain the stylistic features of machine-generated code. Once trained, the models are used to generate new “machine-humanised” code samples. We filter their outputs as in Section 3.6 to keep only high-quality adversarial examples. Finally, we obtained a nearly 1:1 ratio of prompt-based vs. preference-tuning adversarial attacks.
@@ 3.5 Varying Decoding Strategies
It was shown by Ippolito et al. (2020) that after greedy decoding, it was easier to detect AI-generated texts compared to when using other decoding techniques. To capture this challenging behaviour and reflect the diversity of generative systems, we experimented with various decoding strategies, as shown in Table 2.
@@ 3.6 Data Filtering
It should be acknowledged that we did not filter out the human-written codes which were sourced after the coding copilots became popular. It means, that there is a possibility that among human written codes there could be samples, created with the help of LLMs. This potential issue is tackled in Section 5.
@@ 4.1 Experimental Setup
Additionally, we train other models directly on our dataset using a multi-class objective: (i) a simple GCN(Kipf and Welling, 2017); (ii) a CatBoost classifier(Prokhorenkova et al., 2018), following procedures similar to the Whodunit paper; and (iii) two encoder-only transformers, ModernBERT-Base and ModernBERT-Large (Warner et al., 2024), denoted as DroidDetectCLS-Base and DroidDetectCLS-Large. Full details are provided in Sections C.1, C.2 and 5.
@@ 4.2 RQ1: What is the value of extensive data collection for training robust detectors?
Table 1 highlights a key limitation of existing datasets: they often lack diversity. Tables 3 and 4 show that this limitation significantly impacts detector performance in realistic settings. Baseline detectors - the ones illustrated in Zero-Shot Baselines section of the tables - underperform on our test split compared to even simple fine-tuned baselines like GCN and CatBoost. Among the baselines evaluated, zero-shot Fast-DetectGPT consistently yields strong performance across both languages and domains, outperforming all other baselines. In contrast, pre-trained models usually perform well only on languages and domains that are closely aligned with their original training data. This highlights the limitations of previously collected datasets, which do not cover the diversity of generations in DroidCollection, and hence are far from being useful in real-life scenarios.
@@ 4.3 RQ2: How well do models generalise in OOD settings?
Table 4 shows that, under restricted training conditions, models tend to generalise better to syntactically similar languages. For example, a model trained on C/C++ performs reasonably well on C# and Java. However, for typologically isolated languages such as Python or JavaScript, all models not trained specifically for it tend to struggle in this setting. Table 3 illustrates that the models trained on a single domain have a high discrepancy in scores for other domains. For example, as can be seen in the OOD Evaluation section, the classifiers trained on algorithmic problems suffer with the general codes, and show comparable low performance on Research/DS codes, and vice versa.
@@ 4.4 RQ3: How robust are models to adversarial samples?
Finally, we test adversarial robustness using challenging samples designed to evade detection (Table 5). Here, many baseline detectors struggle: for instance, GPT-Zero achieves only 0.10 recall. M4 and CoDet-M4 show higher recall on adversarial samples but also have high false positives, misclassifying human-written code. Interestingly, fine-tuning on DroidCollection also improves the recall of baselines for human-written cases. Remarkably, the recall of M4 on adversarial cases drops from 0.73 to 0.67, while the recall on human-written cases increases significantly from 0.40 to 0.91. A similar pattern is observed for CoDet-M4, while for GPT-Sniffer, both scores increase.
@@ 5 Detector Training and Ablations
As an architectural ablation, we explore whether incorporating the structural representation of code could improve the detector’s performance. Specifically, we trained a 4-layer Graph Convolutional Network over the AST representation of codes to evaluate its ability to distinguish AI-generated from human-written code. The results are shown in Section C.1. We can see that while structural signals are informative, GCNs alone are not sufficient to achieve strong generalisation.
@@
Next, we explored early fusion of textual and structural representations by combining a text encoder with a GCN encoder. For text encoding, we used ModernBERT (Warner et al., 2024), a transformer-based model pre-trained on a mixture of natural language and code. We experimented with both the base (149M parameters) and the large (396M parameters) variants. This model was selected for the inference efficiency (Warner et al., 2024), and suitability for code-related tasks. However, as shown in Section C.3, this fusion strategy yielded only a marginal improvement. Consequently, we decided to use a text-only encoder for the final model.
@@
We then address the issue of class separability, which can arise because adversarial and refined code is similar to human-written code. We explore training our models using triplet loss (Hoffer and Ailon, 2015) in a supervised contrastive (Khosla et al., 2020) setup using the class labels. This metric-learning approach encourages the model to place samples of the same class closer to each other in the embedding space while pushing dissimilar samples apart, and it has been demonstrably effective in other detection scenarios that require high precision (Deng et al., 2022; Li and Li, 2024). We refer to these models as DroidDetectSCL. Table 6 demonstrates the small but consistent performance gain unlocked using metric learning.
@@
Finally, we addressed the problem of noisy and mislabelled training data. Despite extensive data filtering, it is possible that some code samples curated as human-written may have been generated by coding-copilots. The presence of such examples could negatively impact the training of our detector. To address this, we applied MC Dropout (Hasan et al., 2022) to estimate the model uncertainty on the human-written portion of the test set. Specifically, we identified the top 7% most uncertain samples– those for which a pre-trained model exhibited low prediction confidence–and resampled the dataset, removing them from the training set. We then retrained the model on the remaining data, thereby getting rid of the influence of potentially mislabelled or ambiguous samples. This manner of self-bootstrapping datasets has an extensive track record in image (Yalniz et al., 2019; Xie et al., 2020) and text (Wang et al., 2022) representation learning, relying on the tendency of neural networks to understand patterns in clean labels before they overfit to noisy data (Feldman and Zhang, 2020). Incorporating this filtering into our training yielded our final DroidDetect models, which, as shown in Table 6, perform the best across the classification settings.
@@ Appendix A List Of Models Used
Table 7 illustrates that we are using a diverse set of models from 11 model families, combining both instruct and base versions of models. We are also covering a diverse set of sizes: from 2B up to 72B, and use both open-weights and API-based models.
@@ B.2 DroidCollection-Personas creation
To generate DroidCollection-Personas, we started by identifying the main characteristics of a programmer. Our final list contains 9 features: Primary Programming Language, Preferred Frameworks, Field of Work, Code Commenting Style, Error-Proneness, Debugging Strategies, Code Aesthetics, Documentation Habits, Function Length Preference. The possible values for each feature are listed in Table 8.
@@ B.3 Dataset Statistics
In this section, we present key statistics of our dataset and compare them with existing alternatives. As shown in Table 9, our dataset includes a broader class distribution and shows greater diversity in code structure, as reflected by higher AST depth percentiles and longer line lengths. It suggests that our dataset captures more complex and varied code patterns, making it a more challenging and real-life-oriented benchmark for evaluating AI-generated code detection models. The importance of varying code lengths and difficulties is also shown in Section D.1. We also show the number of samples per generator, and programming language (not considering the datasets with <=<=< = 2 languages or generators). Several qualitative examples of samples belonging to different classes in our dataset are shown in LABEL:tab:code_samples_1 and LABEL:tab:code_samples_2.
@@ C.1 GCN Experiments
As shown in Table 11, features based on the textual content of the node yielded the best performance, showing that the semantic information is important in distinguishing between human-written and AI-generated code.
@@ C.2 CatBoost Experiments
These features were used to train CatBoost classifiers with automatically tuned hyperparameters. Figure 1 shows the top unique features ranked by SHAP (SHapley Additive exPlanations) values (Lundberg and Lee, 2017). Interestingly, the most informative features vary across the 2-, 3-, and 4-class classification tasks, suggesting that different granularities of classification are dependent on different aspects of code structure.
@@ C.3 Does Structure-Based Late-Fusion Improve Robustness?
To decide whether fusion is helpful for improving the detection, we combined the GCN from Section C.1 with our text-only classifier using early fusion of embeddings. We used OOD-based generalisation, and compared how well the models perform for 2, 3, and 4-class classification in OOD settings (since when trained directly, it is hard to measure the significance of the performance difference), and then compared in which scenarios each method provides a better weighted F1-score. Table 12 shows that there is no clear trend of one approach being better than another: in the binary classification task, there are more ties, fusion has a higher win-rate in 4-class classification, while the model without fusion performs best in the 3-class case. Then we compared how the difference in F1-scores between models compares to the interquartile range within the model’s predictions. As shown in Figure 2, the interquartile range is much larger than the model difference, so both models with and without fusion perform nearly equally.

### v3

@@ Abstract
Daniil Orel1, Indraneil Paul2, Iryna Gurevych1,2, and Preslav Nakov1 1 Mohamed bin Zayed University of Artificial Intelligence (MBZUAI), UAE 2 Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science TU Darmstadt and National Research Center for Applied Cybersecurity ATHENE, Germany {name.surname}@mbzuai.ac.ae; {name.surname}@tu-darmstadt.de
@@ 2.1 AI-Generated Text Detection
Tools such as MOSS (Puryear and Sprint, 2022) have shown some effectiveness in identifying AI-generated code, since their style is out of the ordinary distributions of student solutions. However, Pan et al. (2024) and JianWang et al. (2024) have shown that detectors such as GPT-Zero often fail when applied to code rather than text. This critical observation, backed up by our experiments in Section˜4, highlights the inadequacy of directly porting generic text-based models to the code domain and strongly motivates the creation of code-specific detection strategies and specialised datasets. Our work responds to this need by providing a large-scale, multifaceted suite specifically curated for AI-generated code, designed to foster the development and rigorous testing of detection techniques attuned to the unique characteristics of programming languages and AI-generated software.
@@ 3 The DroidCollection Corpus
We compare our dataset with existing ones in Table˜1, which shows that our dataset is not only among the largest to date, but also captures a broader range of variations. Additional details about key characteristics are provided in Section˜B.3.
In this section, we detail the curation of the human-generated, machine-generated, and machine-refined splits of DroidCollection. The adversarially humanised data collection is deferred to Section˜3.4.
@@ 3.1 Human-Authored Code Acquisition
To build the dataset, we collected human-written samples from multiple sources, covering C++/C, C#, Go, Java, JavaScript and Python languages. Then, we generated code, using base and instruction-tuned LMs from 11 model release families, namely: Llama (Grattafiori et al., 2024), CodeLlama, GPT-4o, Qwen (Qwen et al., 2025), IBM Granite (Mishra et al., 2024), Yi (AI et al., 2025), DeepSeek (Guo et al., 2024), Phi (Abdin et al., 2024), Gemma (Gemma et al., 2024), Mistral (AI, 2025), Starcoder (Li et al., 2023c). The list of generators per model family is detailed in Appendix˜A. Our dataset covers three domains: general use code, algorithmic problems, and research/data-science code.
@@ 3.2 AI-Authored Code Generation
Since the data from sources such as CodeNet and StarcoderData do not contain any instructions, we decided to apply inverse instructions to transform code from these datasets into instructions, which can be used to prompt LMs. In our case, the method of preparing inverse instructions was similar to that described in InverseCoder (Wu et al., 2025): we passed the code snippets to an LM, asking it to build a summary, and a step-by-step instruction that can be given to an LM to generate a similar code. The main difference between our approach and that of InverseCoder is that we sought to minimise the costs of the generation and did not split the summarisation and instruction generation into separate LM calls. However, in cases where a summary could be extracted from the response but the instruction could not, we used the summary to re-generate the instruction. This experiment with details about the prompts and the models we used is illustrated in Section˜B.1. This type of generation allows us to cover a wide range of prompts, simulating a diversity of user-LM interactions, which is common in the real world.
@@
To explore how large language models can simulate the behaviour profiles of real programmers, we took inspiration from the PersonaHub dataset (Ge et al., 2024). We first generated a diverse set of programmer profiles, and then used an LM to create programming tasks that can typically be performed by programmers of such types. These tasks, along with their corresponding descriptions, termed DroidCollection-Personas, were then used to generate code samples. More details about the DroidCollection-Personas are outlined in the Section˜B.2.
@@ 3.4 Adversarial Samples
To this end, we introduce two types of adversarial attacks: prompt-based attacks and preference-tuning-based attacks. In the prompt-based setting, we construct adversarial prompts by instructing the model to “write like a human” in multiple ways, relying on the models’ parametric knowledge of how to produce outputs that mimic human-authored code and thus challenge detection systems. In the preference-tuning-based setting, we curate DroidCollection-Pref, a dataset of 157K paired examples consisting of human-written and LM-generated code responses to the same prompt. Using DroidCollection-Pref, we train LMs with up to 9B parameters –including LLaMA, Qwen, and Yi–, using LoRA (Hu et al., 2022) with rank 128 and DPO for two epochs. These models’ output distributions are, in effect, steered towards preferring human-like code, making them less likely to contain the stylistic features of machine-generated code. Once trained, the models are used to generate new “machine-humanised” code samples. We filter their outputs as in Section˜3.6 to keep only high-quality adversarial examples. Finally, we obtained a nearly 1:1 ratio of prompt-based vs. preference-tuning adversarial attacks.
@@ 3.5 Varying Decoding Strategies
It was shown by Ippolito et al. (2020) that after greedy decoding, it was easier to detect AI-generated texts compared to when using other decoding techniques. To capture this challenging behaviour and reflect the diversity of generative systems, we experimented with various decoding strategies, as shown in Table˜2.
@@ 3.6 Data Filtering
It should be acknowledged that we did not filter out the human-written codes which were sourced after the coding copilots became popular. It means, that there is a possibility that among human written codes there could be samples, created with the help of LLMs. This potential issue is tackled in Section˜5.
@@ 4.1 Experimental Setup
Additionally, we train other models directly on our dataset using a multi-class objective: (i) a simple GCN(Kipf and Welling, 2017); (ii) a CatBoost classifier(Prokhorenkova et al., 2018), following procedures similar to the Whodunit paper; and (iii) two encoder-only transformers, ModernBERT-Base and ModernBERT-Large (Warner et al., 2024), denoted as DroidDetectCLS-Base and DroidDetectCLS-Large. Full details are provided in Sections˜C.1, C.2 and 5.
@@ 4.2 RQ1: What is the value of extensive data collection for training robust detectors?
Table˜1 highlights a key limitation of existing datasets: they often lack diversity. Tables˜3 and 4 show that this limitation significantly impacts detector performance in realistic settings. Baseline detectors - the ones illustrated in Zero-Shot Baselines section of the tables - underperform on our test split compared to even simple fine-tuned baselines like GCN and CatBoost. Among the baselines evaluated, zero-shot Fast-DetectGPT consistently yields strong performance across both languages and domains, outperforming all other baselines. In contrast, pre-trained models usually perform well only on languages and domains that are closely aligned with their original training data. This highlights the limitations of previously collected datasets, which do not cover the diversity of generations in DroidCollection, and hence are far from being useful in real-life scenarios.
@@ 4.3 RQ2: How well do models generalise in OOD settings?
Table˜4 shows that, under restricted training conditions, models tend to generalise better to syntactically similar languages. For example, a model trained on C/C++ performs reasonably well on C# and Java. However, for typologically isolated languages such as Python or JavaScript, all models not trained specifically for it tend to struggle in this setting. Table˜3 illustrates that the models trained on a single domain have a high discrepancy in scores for other domains. For example, as can be seen in the OOD Evaluation section, the classifiers trained on algorithmic problems suffer with the general codes, and show comparable low performance on Research/DS codes, and vice versa.
@@ 4.4 RQ3: How robust are models to adversarial samples?
Finally, we test adversarial robustness using challenging samples designed to evade detection (Table˜5). Here, many baseline detectors struggle: for instance, GPT-Zero achieves only 0.10 recall. M4 and CoDet-M4 show higher recall on adversarial samples but also have high false positives, misclassifying human-written code. Interestingly, fine-tuning on DroidCollection also improves the recall of baselines for human-written cases. Remarkably, the recall of M4 on adversarial cases drops from 0.73 to 0.67, while the recall on human-written cases increases significantly from 0.40 to 0.91. A similar pattern is observed for CoDet-M4, while for GPT-Sniffer, both scores increase.
@@ 5 Detector Training and Ablations
As an architectural ablation, we explore whether incorporating the structural representation of code could improve the detector’s performance. Specifically, we trained a 4-layer Graph Convolutional Network over the AST representation of codes to evaluate its ability to distinguish AI-generated from human-written code. The results are shown in Section˜C.1. We can see that while structural signals are informative, GCNs alone are not sufficient to achieve strong generalisation.
@@
Next, we explored early fusion of textual and structural representations by combining a text encoder with a GCN encoder. For text encoding, we used ModernBERT (Warner et al., 2024), a transformer-based model pre-trained on a mixture of natural language and code. We experimented with both the base (149M parameters) and the large (396M parameters) variants. This model was selected for the inference efficiency (Warner et al., 2024), and suitability for code-related tasks. However, as shown in Section˜C.3, this fusion strategy yielded only a marginal improvement. Consequently, we decided to use a text-only encoder for the final model.
@@
We then address the issue of class separability, which can arise because adversarial and refined code is similar to human-written code. We explore training our models using triplet loss (Hoffer and Ailon, 2015) in a supervised contrastive (Khosla et al., 2020) setup using the class labels. This metric-learning approach encourages the model to place samples of the same class closer to each other in the embedding space while pushing dissimilar samples apart, and it has been demonstrably effective in other detection scenarios that require high precision (Deng et al., 2022; Li and Li, 2024). We refer to these models as DroidDetectSCL. Table˜6 demonstrates the small but consistent performance gain unlocked using metric learning.
@@
Finally, we addressed the problem of noisy and mislabelled training data. Despite extensive data filtering, it is possible that some code samples curated as human-written may have been generated by coding-copilots. The presence of such examples could negatively impact the training of our detector. To address this, we applied MC Dropout (Hasan et al., 2022) to estimate the model uncertainty on the human-written portion of the test set. Specifically, we identified the top 7% most uncertain samples– those for which a pre-trained model exhibited low prediction confidence–and resampled the dataset, removing them from the training set. We then retrained the model on the remaining data, thereby getting rid of the influence of potentially mislabelled or ambiguous samples. This manner of self-bootstrapping datasets has an extensive track record in image (Yalniz et al., 2019; Xie et al., 2020) and text (Wang et al., 2022) representation learning, relying on the tendency of neural networks to understand patterns in clean labels before they overfit to noisy data (Feldman and Zhang, 2020). Incorporating this filtering into our training yielded our final DroidDetect models, which, as shown in Table˜6, perform the best across the classification settings.
@@ 6 Conclusion and Future Work
Acknowledgements
At TU Darmstadt, this research work has been supported by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE and by HUAWEI Technologies (Ireland) Co., Ltd.
@@ Appendix A List Of Models Used
Table˜7 illustrates that we are using a diverse set of models from 11 model families, combining both instruct and base versions of models. We are also covering a diverse set of sizes: from 2B up to 72B, and use both open-weights and API-based models.
@@ B.2 DroidCollection-Personas creation
To generate DroidCollection-Personas, we started by identifying the main characteristics of a programmer. Our final list contains 9 features: Primary Programming Language, Preferred Frameworks, Field of Work, Code Commenting Style, Error-Proneness, Debugging Strategies, Code Aesthetics, Documentation Habits, Function Length Preference. The possible values for each feature are listed in Table˜8.
@@ B.3 Dataset Statistics
In this section, we present key statistics of our dataset and compare them with existing alternatives. As shown in Table˜9, our dataset includes a broader class distribution and shows greater diversity in code structure, as reflected by higher AST depth percentiles and longer line lengths. It suggests that our dataset captures more complex and varied code patterns, making it a more challenging and real-life-oriented benchmark for evaluating AI-generated code detection models. The importance of varying code lengths and difficulties is also shown in Section˜D.1. We also show the number of samples per generator, and programming language (not considering the datasets with <=<= 2 languages or generators). Several qualitative examples of samples belonging to different classes in our dataset are shown in LABEL:tab:code_samples_1 and LABEL:tab:code_samples_2.
@@ C.1 GCN Experiments
As shown in Table˜11, features based on the textual content of the node yielded the best performance, showing that the semantic information is important in distinguishing between human-written and AI-generated code.
@@ C.2 CatBoost Experiments
These features were used to train CatBoost classifiers with automatically tuned hyperparameters. Figure˜1 shows the top unique features ranked by SHAP (SHapley Additive exPlanations) values (Lundberg and Lee, 2017). Interestingly, the most informative features vary across the 2-, 3-, and 4-class classification tasks, suggesting that different granularities of classification are dependent on different aspects of code structure.
@@ C.3 Does Structure-Based Late-Fusion Improve Robustness?
To decide whether fusion is helpful for improving the detection, we combined the GCN from Section˜C.1 with our text-only classifier using early fusion of embeddings. We used OOD-based generalisation, and compared how well the models perform for 2, 3, and 4-class classification in OOD settings (since when trained directly, it is hard to measure the significance of the performance difference), and then compared in which scenarios each method provides a better weighted F1-score. Table˜12 shows that there is no clear trend of one approach being better than another: in the binary classification task, there are more ties, fusion has a higher win-rate in 4-class classification, while the model without fusion performs best in the 3-class case. Then we compared how the difference in F1-scores between models compares to the interquartile range within the model’s predictions. As shown in Figure˜2, the interquartile range is much larger than the model difference, so both models with and without fusion perform nearly equally.